I'm suppose to solve the system of equation by using the addition method:
x-y =-9
x+3 = 15
I know the answer is (-3,6) but I dont know HOW they got the answer.

Answer by Alan3354(69443)

x-y =-9
x+3 = 15
====================
I think you mean
x- y =-9
x+3y = 15
-------------- Subtract
-4y = -24
y = 6
------
Sub for y in either eqn, find x